Sol Summary – May 8th

Crew 261 Sol Summary Report 08-05-2023

Sol 8
Summary Title: Gypsum crystals on the Martian soil
Author’s name: Aline Decadi, XO

Mission Status: nominal

Sol Activity Summary: Today took place 2 EVAs:
– EVA-11 with the following 4 crew members: James Burk (Commander) (EVA Lead), Julien Villa-Massone (Crew Engineer), Cecile Renaud (GreenHab Officer), Kris Davidson (Journalist) (Not In Sim – Not Wearing Suit)
– EVA-12 with the following 4 crew members: Audrey Derobertmasure (HSO), Erin Kennedy (Crew Roboticist), Aline Decadi (XO) (EVA Lead), Kris Davidson (Journalist) (Not In Sim – Not Wearing Suit)

EVA-11 conducted testing of ADAPA 360 Drone. The crew walk outside of Hab to a suitable flat piece of ground for drone launch & landing. This preliminary testing of the drone was successfully, and the EVA was also successful. Another EVA is foreseen to take 3D imaging to map a terrain on the field.

EVA-12 had the following purpose:
1) Find, investigate, and collect Gypsum crystals on the Glistening Seas region from selenite, the transparent form of gypsum. This relates to Atmosphinder experiment as it could be an analogous feature to CO2 gas jet expulsions in the Mars South Polar region.
2) Atmosphinder electronics environmental sensor data logging.
EVA-12 ventured North of the Hab to the vicinity of the Valles Marineris region, located East of Cow Dung Road, and South of Brahe Highway.
Upon approaching the destination, it became evident why this location is nicknamed “Glistening Seas”. Gazing across the horizon, sparkles could be seen from the sunlight reflecting on the crystals. Gypsum crystals protruded from the red-tinged dirt below. The EVA crew investigated a cluster of these and collected samples. The EVA crew investigated this area thoroughly, finding rocks of interest and additional crystals. The crew was elated with the plethora of crystals and fascinating geological features!
Atmosphinder was extracted from the rover for a test. The wind gusts were powerful in this region as it was a clear plain. Atmosphinder was given an initial push, and it traveled ~70 m as propelled by the wind. Additional testing occurred, with the same results. The result was compelling to observe a structure moving on its own, venturing to explore the vastness of Mars! After a test concluded in disassembly, the area was thoroughly checked for remaining pieces, and the pieces were brought back to the rover.
Crew health checks were completed regularly by the EVA lead, everyone was feeling great and in great spirits. We received coordinates to a point of interest of a crew member’s lost drone. We conducted a search for the lost drone, akin to a ‘geocaching’ activity. The search concluded without retrieving the drone, though admittedly, admiring the incredible geology at that point of interest.
The EVA-12 resulted in awe-inspiring observations and making connections of those observations to Mars. New information was learned about Atmosphinder robot that would only be possible in this unique environment. There was an experiential confidence boost for some of the EVA crew. The collected crystal samples will be further investigated through a microscope, shared, and whimsically gazed upon. Many thanks to the entire crew — both on the EVA, at HabCom, and behind the scenes!

The crew spent the rest of the day: Audrey Derobertmasure (HSO) performed a demonstration on her thesis on Cardiovascular measurements and the rigorous protocol under development for handling and storing human biological samples (as part of the COSMOS experiment); having dinner together and preparing the next day’s experiments and EVAs.

Look Ahead Plan: we will conduct 2 EVAs tomorrow: EVA-13 in the morning for the rescue of Julien’s drone, and fly ADAPA drone to get footage; EVA-14 in the evening for photoshoot in the immediate vicinity of the Hab, according to mission request approval.

Anomalies in work: None.

Weather: Pleasant and sunny. Very windy in the afternoon.

Crew Physical Status: Nominal.
EVA: EVA-11 in the vicinity of the Hab; EVA-12 on the North of Cow Dung Road.

Reports to be filed: HSO Report, EVA Report, Operations Report, Journalist Report, Pictures of the day, GreenHab Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site